A farmer has a 10% insecticide solution and a 50% solution. How much of each should he mix together to obtain 40 gallons of a 35% solution?
Answer by checkley77(12569)

.5X+.1(40-X)=.35*40
.5X+4-.1X=14
.4X=14-4
.4X=10
X=10/.4
X=25 GAL. OF 50% INSECTICIDE IS USED.
40-25=15 GAL. OF 10% INSECTICIDE IS USED.
PROOF:
.5*25+.1*15=.35840
12.5+1.5=14
14=14